### v3

@@ Abstract
The emergence of multimodal large language models has redefined the agent paradigm by integrating language and vision modalities with external data sources, enabling agents to better interpret human instructions and execute increasingly complex tasks. However, in this work, we identify a critical yet previously overlooked security vulnerability in multimodal agents: cross-modal prompt injection attacks. To exploit this vulnerability, we propose CrossInject, a novel attack framework in which attackers embed adversarial perturbations across multiple modalities to align with target malicious content, allowing external instructions to hijack the agent’s decision-making process and execute unauthorized tasks. Our approach consists of two key components. First, we introduce Visual Latent Alignment, where we optimize adversarial features to the malicious instructions in the visual embedding space based on a text-to-image generative model, ensuring that adversarial images subtly encode cues for malicious task execution. Subsequently, we present Textual Guidance Enhancement, where a large language model is leveraged to infer the black-box defensive system prompt through adversarial meta prompting and generate an malicious textual command that steers the agent’s output toward better compliance with attackers’ requests. Extensive experiments demonstrate that our method outperforms existing injection attacks, achieving at least a +26.4% increase in attack success rates across diverse tasks. Furthermore, we validate our attack’s effectiveness in real-world multimodal autonomous agents, highlighting its potential implications for safety-critical applications.
@@ 1 Introduction
he rapid advancement of large vision language models (LVLMs) Xie et al. (2024); Liu et al. (2023, 2024); Dai et al. (2023); Zhuang et al. (2025); Chen et al. (2024); Yao et al. (2024); Wang et al. (2024) has significantly enhanced the capabilities of multimodal agents, enabling them to process and integrate heterogeneous information sources such as visual scenes, natural language instructions, and structured external data. These multimodal agents have demonstrated impressive performance across a wide range of applications, including virtual assistants Pei et al. (2024); Chu et al. (2023); Hong et al. (2024), autonomous driving systems Wang et al. (2023); Ma et al. (2024); Xu et al. (2024), and embodied agents Abdin et al. (2024); Wallace et al. (2024); Mu et al. (2023); Li et al. (2024); Qin et al. (2023). By leveraging the complementary strengths of different modalities, these systems can make more informed and context-sensitive decisions.
@@
However, the increasing reliance on user-provided and external inputs introduces new security vulnerabilities known as prompt injection attacks. By embedding adversarial content within the agent’s input streams, the adversaries can manipulate the behavior of an agent to execute unauthorized tasks OWASP (2023). While such attacks have been extensively studied in unimodal settings (targeting either the visual Shayegani et al. (2024); Wu et al. (2024) or textual Liu et al. (2024, 2023) modality), these attacks fail to account for the complex interactions between modalities that fundamentally drive the behavior of multimodal agents. However, it is highly non-trivial to perform effective prompt injection attacks in the multimodal setting. Unlike unimodal systems, multimodal agents integrate semantically rich inputs from diverse modalities, which collectively influence the agent’s decision-making process. As a result, attacks must be carefully coordinated across modalities to exploit the vulnerabilities introduced by multimodal data fusion. Despite the growing deployment of such agents, there remains a significant gap in the literature: there is no systematic attack framework tailored to black-box multimodal agents, nor a comprehensive understanding of the risks associated with cross-modal interactions.
@@
To address this gap, this paper introduces the concept of cross-modal prompt injection and proposes CrossInject, a novel prompt injection framework specifically designed for black-box multimodal agents. In contrast to conventional attacks that target a single modality or input channel (such as user prompt Liu et al. (2023) or external data Abdelnabi et al. (2023)), CrossInject introduces a coordinated cross-modal attack strategy. This strategy simultaneously injects adversarial contents into both visual and textual modalities across multiple input sources, fully exploiting the vulnerabilities introduced by multimodal data fusion. CrossInject  introduces a cross-modal prompt injection strategy that targets both visual and textual modalities through two complementary components. The first component, Visual Latent Alignment, leverages a text-to-image generative model to synthesize images that are semantically aligned with malicious instructions. These generated images are embedded within benign visual contexts, enabling the injection of adversarial semantics into the visual modality while preserving plausibility and stealth. To further enhance control over the agent’s behavior, CrossInject incorporates Textual Guidance Enhancement, which optimizes malicious textual commands using surrogate open-source large language model. This component is designed to steer the agent towards the attacker-specified goals through textual manipulation. Combined with malicious instructions embedded in external data sources, which is a known vulnerability in real-world agent deployments Abdelnabi et al. (2023); Yi et al. (2025); Zhan et al. (2024), these components form a highly effective cross-modal prompt injection pipeline targeting multiple modalities from various input sources.
@@
We conduct extensive experiments across diverse multimodal agent architectures and real-world scenarios. Our results show that CrossInject consistently outperforms existing prompt injection methods, achieving at least 26.4% higher average attack success rates. In addition, CrossInject demonstrates strong transferability across different agent models, confirming its effectiveness in black-box and cross-modal settings. We further evaluate CrossInject against representative defense strategies, and the results indicate that these defenses are largely ineffective against our attacks. Furthermore, we validate our attack’s effectiveness on real-world multimodal autonomous agents, highlighting its potential implications in practice. Our contributions are summarized as follows:
@@
  • We conduct comprehensive empirical evaluations across multiple multimodal agent architectures over different tasks in both digital and physical worlds. CrossInject achieves +26.4% attack success rates compared to existing prompt injection methods.
@@ 2.1 Multimodal Agent
As multimodal agents gain widespread adoption across critical domains, significant security vulnerabilities have emerged Liu et al. (2024); Wang et al. (2025); Zhang et al. (2025); Wu et al. (2024); Aichberger et al. (2025); Liu et al. (2020); Liang et al. (2020); Wei et al. (2018); Liang et al. (2022a, b); Yan et al. (2024); Ying and Wu (2023a, b). In this paper, we investigate the security implications Zhang et al. (2024); Kong et al. (2024); ying2024safebench; Ying et al. (2025); Liang et al. (2023, 2024a, 2024b); Ying et al. (2024a, b, 2025); Jing et al. (2025) of multimodal agents under cross-modal prompt injection attacks, with particular emphasis on their susceptibility to malicious inputs originating from diverse sources.
@@ 3.1 Problem Definition
Cross-Modal Prompt Injection. Conventional prompt injection attacks mainly apply adversarial operation on unimodal input. However, in the context of cross-modal prompt injection against multimodal agents, the agent’s input consists of multiple sources from distinct modalities, requiring coordinated perturbations across various data streams. Specifically, we design a visual injection function $\phi(\cdot)$, which transforms the $I$ into $I+\delta_{I}$ with the adversarial perturbation $\delta_{I}$, and textual injection function $\eta(\cdot)$, which transforms $C$ into $C^{\prime}$ embedding deceptively semantics that manipulates reasoning outcomes. In addition, to exploit the vulnerability introduced by external data interfaces, we design malicious external data $\psi(E,d)$, where $\psi(\cdot,\cdot)$ is a transform that injects malicious instruction $d$ into the external data $E$, generating manipulated data $E^{\prime}$. After feeding the whole adversarial input into the victim agent, the malicious action $a^{*}$ generated by the agent is defined as below:
$\begin{split}a^{*}&=A(S,\ \phi(I),\ \psi(E,d),\ \eta(C))\\ &=A(S,\ I+\delta_{I},\ E^{\prime},\ C^{\prime}).\end{split}$ (2)
@@ 4 Methodology
In this section, we introduce CrossInject, a novel prompt injection attack targeting multimodal agents. Our approach exploits vulnerabilities across multiple modalities, manipulating both visual and textual inputs to steer the agent toward executing malicious tasks. The overall framework is illustrated in Fig. 2.
@@ 4.1 Visual Latent Alignment
Multimodal agents built on VLMs exhibit greater susceptibility to prompt injection attacks compared to unimodal LLMs, due to the additional vulnerabilities introduced by the visual modality with high-dimensional feature space. While existing prompt injection methods predominantly focus on unimodal textual attacks, we propose exploiting the visual modality by injecting malicious content into the visual input $I$
@@
A straightforward approach would be to directly align malicious textual instructions with input images in the latent feature space, embedding malicious semantics within visually benign input Cui et al. (2024). However, achieving cross-modal semantic alignment is challenging due to the inherent modality gap, where textual semantics and visual features often reside in distinct latent spaces with limited correspondence Liang et al. (2022). To address this, inspired by Zhao et al. (2023), we avoid direct alignment of malicious instructions with visual features. Instead, we leverage a text-to-image model $g(\cdot)$ (e.g., Stable Diffusion Rombach et al. (2022)) to generate target images that inherently encode the semantics of the malicious instructions into visual inputs.
@@
Specifically, let the malicious instruction be denoted as $d$ (e.g.. Help me polish my words.). We reformulate $d$ into a text-to-image descriptive prompt $d^{\prime}$ (i.e., A person is polishing his words in an empty room.) to keep semantic relevance between the generated images and the malicious instructions through three key modifications. This transformation successfully grounds textual commands into visual semantics by specifying concrete subject to establish action entity, embedding environmental context for spatial coherence, and employing progressive tense for temporal scene dynamics, thereby enabling precise cross-modal semantic mapping. The target image is then generated as $I_{t}=g(d^{\prime})$. Next, we perform visual feature alignment between the benign input image $I$ and the generated image $I_{t}$ to fully embed malicious content into $I$. We also validate the attack efficacy of directly aligning with textual instruction in Sec. 5.3.
@@
Given our black-box threat model, we adopt a transfer attack strategy, using publicly available pretrained vision encoders (e.g., CLIP Radford et al. (2021), SigLIP Zhai et al. (2023)) as surrogate models to craft adversarial visual input embedded with malicious content. To enhance attack transferability to unseen victim agents, we employ a model ensemble attack strategy, simultaneously attacking multiple surrogate models $\{f_{k}(\cdot)\}_{k=1}^{K}$, where $K$ is the number of surrogate models. These surrogate models are selected based on their widespread adoption in multimodal systems with diverse architectural representations. The loss function for visual latent alignment, $\mathcal{L}_{v}$, is defined as:
@@
$\begin{split}\mathcal{L}_{v}(\delta_{I})&=\frac{1}{K}\sum_{k=1}^{K}\left\|\frac{f_{k}(I+\delta_{I})}{\left\|f_{k}(I+\delta_{I})\right\|_{2}}-\frac{f_{k}(I_{t})}{\left\|f_{k}(I_{t})\right\|_{2}}\right\|_{2},\\ \\ \text{s.t.}\ &\left\|\ \delta_{I}\ \right\|_{\infty}\leq\epsilon.\end{split}$ (3)
@@
The perturbation $\delta_{I}$ is constrained within an $\epsilon$-ball to balance attack effectiveness and imperceptibility, ensuring the $I+\delta_{I}$ remains visually indistinguishable from $I$. We minimize the $\ell_{\infty}$-norm distance between normalized feature representations to ensure scale-invariant alignment in the visual latent space, enhancing robustness of optimization process across different vision encoders.
@@
Specifically, we optimize Eq. (3) using the SSA-CWA Chen et al. (2024) algorithm, which iteratively computes gradients across the ensemble of surrogate models to produce the visual perturbation $\delta_{I}$, thereby generating adversarial images embedded with malicious contents.
@@ 4.2 Textual Guidance Enhancement
While the Visual Latent Alignment manipulates the agent’s visual input to prioritize malicious instructions, it alone may not guarantee precise control over the agent’s output. To overcome this, we inject malicious content into the textual command input $C$ using injection function $\eta(\cdot)$ defined in Sec. 3.1, converting it into a deceptive command $C^{\prime}$. This manipulated command works synergistically with the malicious instruction $d$ embedded in visual input $I+\delta_{I}$ and external data $E^{\prime}$, steering the agent’s output towards compliance with attacker’s request. This forms a cross-modal bidirectional attack strategy, fully exploiting vulnerabilities across the entire input-output pipeline of multimodal agents.
@@
Specifically, we initialize $C^{\prime}$ from random string and iteratively optimize it to maximize the likelihood of generating malicious action $a^{*}$ defined in Eq. (2), inducing the agent to generate output aligned with attacker’s adversarial goal. However, directly optimizing this objective with respect to $C^{\prime}$ under black-box setting is challenging, as multimodal agents typically have high-dimensional inputs and complex multimodal data processing pipelines, making end-to-end gradient computation infeasible Zhang et al. (2025). To address this, we adopt a transfer attack strategy, leveraging open-source LLM as surrogate model to approximate the victim agent’s inference process. We optimize the adversarial textual command $C^{\prime}$ by aligning the surrogate model’s output with a target sequence that reflects the attacker-desired malicious behavior. Let the attacker-desired content be denoted as $T$, the loss function for textual optimization, $\mathcal{L}_{t}$, is defined as:
$\mathcal{L}_{t}(C^{\prime})=-\log p(T\mid S,\ d,\ C^{\prime}),$ (4)
where we minimize the negative log likelihood of the target content $T$ to maximize the probability of the agent generating response aligned with the malicious instruction $d$. In the context of prompt injection attacks, $T$ represents affirmative response prefix that mimics the agent’s potential reply pattern when complying with the attacker’s requests. Specifically, to optimize Eq. (4), we employ the GCG algorithm, which has proven effective in transfer attacks Zou et al. (2023). Since the configuration of external data is dynamic and flexible, we choose to optimize Eq. (4) based on pure malicious instruction $d$ instead of the whole manipulated external data $E^{\prime}$. The impact of surrogate LLM selection on textual injection success rates is quantified in Sec. 5.3.
@@
Due to the attacker’s limited black-box access to the agent system, we consider inferring the defense-aware system prompt $S$ aligned with the agent’s safety mechanism. Here we employ a powerful language model (i.e., GPT-4 OpenAI (2023)) to infer safety-aligned system prompt through role-driven analysis. Typically, to amplify adversarial impact of textual injection, we assume strengthened defensive protection of the victim agent when inferring system prompt, pushing the textual optimization to discover the command $C^{\prime}$ that overcomes such strong protection. This adversarially leads to malicious command $C^{\prime}$ with enhanced control over the victim agent’s output. Specifically, we apply meta prompting Zhang et al. (2025) method to construct the inference prompts for different agents. Meta prompt is an example-agnostic structured prompt, which can be used to efficiently guide LLM to infer system prompts for agents with diverse roles. Let $R$ be denoted as role description of a victim agent, which is known to the attacker in our threat model, and let $D(\cdot)$ represent the defense-aware meta prompt template, the inferred system prompt $S^{\prime}$ is formulated as below:
$S^{\prime}=\mathcal{M}(D(R)),$ (5)
where $\mathcal{M}$ indicates the LLM for system prompts inference. The inferred $S^{\prime}$ is used to replace $S$ in Eq. (4). We also validate the attack efficacy of textual optimization on real system prompt in Sec. 5.3. Detailed inference prompts are provided in supplementary material.

### v4

@@ Abstract
The emergence of multimodal large language models has redefined the agent paradigm by integrating language and vision modalities with external data sources, enabling agents to better interpret human instructions and execute increasingly complex tasks. However, in this paper, we identify a critical yet previously overlooked security vulnerability in multimodal agents: cross-modal prompt injection attacks. To exploit this vulnerability, we propose CrossInject, a novel attack framework in which attackers embed adversarial perturbations across multiple modalities to align with target malicious content, allowing external instructions to hijack the agent’s decision-making process and execute unauthorized tasks. Our approach incorporates two key coordinated components. First, we introduce Visual Latent Alignment, where we optimize adversarial features to the malicious instructions in the visual embedding space based on a text-to-image generative model, ensuring that adversarial images subtly encode cues for malicious task execution. Subsequently, we present Textual Guidance Enhancement, where a large language model is leveraged to construct the black-box defensive system prompt through adversarial meta prompting and generate an malicious textual command that steers the agent’s output toward better compliance with attackers’ requests. Extensive experiments demonstrate that our method outperforms state-of-the-art attacks, achieving at least a +30.1% increase in attack success rates across diverse tasks. Furthermore, we validate our attack’s effectiveness in real-world multimodal autonomous agents, highlighting its potential implications for safety-critical applications.
@@ 1 Introduction
The rapid advancement of large vision language models (LVLMs) Xie et al. (2024); Liu et al. (2023, 2024); Dai et al. (2023); Chen et al. (2024); Yao et al. (2024); Wang et al. (2024); Zhuang et al. (2025) has significantly enhanced the capabilities of multimodal agents, enabling them to process and integrate heterogeneous information sources such as visual scenes, natural language instructions, and structured external data. These multimodal agents have demonstrated impressive performance across a wide range of applications, including virtual assistants Pei et al. (2024); Chu et al. (2023), autonomous driving systems Wang et al. (2023); Ma et al. (2024), and embodied agents Abdin et al. (2024); Wallace et al. (2024). By leveraging the complementary strengths of different modalities, these systems can make more informed and context-sensitive decisions.
@@
However, the increasing reliance on user-provided external inputs introduces new security vulnerabilities known as prompt injection attacks. By embedding adversarial content within the agent’s input streams, the attackers can manipulate the behavior of an agent to execute unauthorized tasks OWASP (2023), causing resource abuse. While such attacks have been extensively studied in unimodal settings (targeting either the visual Shayegani et al. (2024); Wu et al. (2024) or textual Liu et al. (2024, 2023) modality), these attacks fail to account for the complex interactions between modalities that fundamentally drive the behavior of multimodal agents. However, it is highly non-trivial to perform effective prompt injection attacks in the multimodal setting. Unlike unimodal systems, multimodal agents integrate semantically rich inputs from diverse modalities, which collectively influence the agent’s decision-making process. As a result, attacks must be carefully coordinated across modalities to exploit the vulnerabilities introduced by multimodal data fusion. Despite the growing deployment of such multimodal systems, there remains a significant gap in the literature: there is no systematic attack framework tailored to black-box multimodal agents, nor a comprehensive understanding of the risks associated with cross-modal interactions.
@@
To address this gap, this paper introduces the concept of cross-modal prompt injection and proposes CrossInject, a novel prompt injection framework specifically designed for black-box multimodal agents. In contrast to conventional attacks that target a single modality or input channel, CrossInject introduces a coordinated cross-modal attack strategy. This strategy simultaneously injects adversarial contents into both visual and textual modalities across multiple input sources, fully exploiting the vulnerabilities introduced by multimodal data fusion. CrossInject attacks both visual and textual modalities through two complementary components. The first component, Visual Latent Alignment, leverages a text-to-image generative model to synthesize images that are semantically aligned with malicious instructions. These generated images are embedded within benign visual contexts, enabling the injection of adversarial semantics into the visual modality while preserving plausibility and stealth. To further enhance control over the agent’s behavior, CrossInject incorporates Textual Guidance Enhancement, which optimizes malicious textual commands using a surrogate open-source large language model. This component is designed to steer the agent towards the attacker-specified goals through textual manipulation. Combined with malicious instructions embedded in external data sources (a known vulnerability in real-world agent deployments Abdelnabi et al. (2023); Yi et al. (2025); Zhan et al. (2024)), these components form a highly effective cross-modal prompt injection pipeline.
@@
We conduct extensive experiments across diverse multimodal agent architectures and real-world scenarios. Our results show that CrossInject consistently outperforms existing prompt injection methods, achieving at least 30.1% higher average attack success rates. In addition, CrossInject demonstrates strong transferability across different agent models, confirming its effectiveness in black-box settings. We further evaluate CrossInject against representative defense strategies, and the results indicate that these defenses are ineffective against our attack. Furthermore, we validate our attack’s effectiveness in real-world multimodal autonomous driving agents, highlighting its potential implications in practice. Our contributions are summarized as follows:
@@
  • We conduct comprehensive empirical evaluations across multiple multimodal agents over different tasks in both digital and physical worlds. CrossInject achieves +30.1% attack success rates compared to existing prompt injection methods.
@@ 2.1 Multimodal Agent
As multimodal agents gain widespread adoption across critical domains, significant security vulnerabilities have emerged Liu et al. (2024); Wang et al. (2025); Zhang et al. (2025); Wu et al. (2024); Aichberger et al. (2025); Liu et al. (2020); Liang et al. (2020); Wei et al. (2018); Liang et al. (2022a, b); Yan et al. (2024). In this paper, we investigate the security implications Zhang et al. (2024); Kong et al. (2024); Ying et al. (2025); Liang et al. (2023, 2024a, 2024b); Ying et al. (2024a, b, 2025); Jing et al. (2025) of multimodal agents under cross-modal prompt injection attacks, with particular emphasis on their susceptibility to malicious inputs originating from diverse sources.
@@ 3.1 Problem Definition
Cross-Modal Prompt Injection. Conventional prompt injection attacks mainly apply adversarial operation on unimodal input. However, in the context of cross-modal prompt injection against multimodal agents, the agent’s input consists of multiple sources from distinct modalities, requiring coordinated perturbations across various data streams. Specifically, we design a visual injection function $\phi(\cdot)$, which transforms the $I$ into $I+\delta$ with the adversarial perturbation $\delta$, and textual injection function $\eta(\cdot)$, which transforms $C$ into $C^{\prime}$ embedding deceptively semantics that manipulates reasoning outcomes. In addition, to exploit the vulnerability introduced by external data interfaces, we design malicious external data $\psi(E,d)$, where $\psi(\cdot,\cdot)$ is a transform that injects malicious instruction $d$ into the external data $E$, generating manipulated data $E^{\prime}$. After feeding the whole adversarial input into the victim agent, the malicious action $a^{*}$ generated by the agent is defined as below:
$\begin{split}a^{*}&=A(S,\ \phi(I),\ \psi(E,d),\ \eta(C))\\ &=A(S,\ I+\delta,\ E^{\prime},\ C^{\prime}).\end{split}$ (2)
@@ 4 Methodology
In this section, we introduce CrossInject, a novel prompt injection attack targeting multimodal agents. Our approach exploits vulnerabilities across multiple modalities, manipulating both visual and textual inputs to steer the agent toward executing injected tasks. The overall framework is illustrated in Fig. 2.
@@ 4.1 Visual Latent Alignment
Multimodal agents built on VLMs exhibit greater susceptibility to prompt injection attacks compared to unimodal LLMs, due to the additional vulnerabilities introduced by the visual modality with high-dimensional feature space. While existing prompt injection methods predominantly focus on unimodal textual attacks, we propose exploiting the visual modality by injecting malicious cues into the visual input $I$.
@@
A straightforward approach would be directly aligning malicious textual instructions with input images in the latent feature space, embedding malicious semantics within visually benign input Cui et al. (2024). However, achieving cross-modal semantic alignment is challenging due to the inherent modality gap, where textual semantics and visual features often reside in distinct latent spaces with limited correspondence Liang et al. (2022). To address this, inspired by Zhao et al. (2023), we avoid direct alignment of malicious instructions with visual features. Instead, we leverage a text-to-image model $g(\cdot)$ to generate target images that inherently encode the semantics of the malicious instructions into visual inputs.
@@
Specifically, given an attacker’s instruction $d$ (e.g., Help me polish my paragraph), we reformulate $d$ into a text-to-image descriptive prompt $d^{\prime}$ (e.g., A person is polishing his paragraph on a table). This reformulation grounds the textual command into visual semantics by specifying subject, context and action for task execution, thereby enabling precise cross-modal semantic mapping from malicious command to visual representation. The target image is then generated as $I_{t}=g(d^{\prime})$ that visually embodies the malicious intent.We then perform visual feature alignment between the benign input image $I$ and the generated image $I_{t}$ to fully embed injected task into $I$.
@@
Given our black-box threat model, we adopt a transfer-based attack strategy. To maximize attack transferability to unseen agents, we craft the adversarial visual input by leveraging an ensemble of $K$ publicly available, pre-trained vision encoders $\{f_{k}(\cdot)\}_{k=1}^{K}$ as surrogate models including CLIP Radford et al. (2021) and SigLIP Zhai et al. (2023). These models are selected based on their widespread adoption in diverse multimodal systems. The loss function for visual injection, $\mathcal{L}_{v}$, is defined as:
@@
$\begin{split}\mathcal{L}_{v}(\delta)&=\frac{1}{K}\sum_{k=1}^{K}\left\|\frac{f_{k}(I+\delta)}{\left\|f_{k}(I+\delta)\right\|_{2}}-\frac{f_{k}(I_{t})}{\left\|f_{k}(I_{t})\right\|_{2}}\right\|_{2},\\ \\ \text{s.t.}\ &\left\|\ \delta\ \right\|_{\infty}\leq\epsilon.\end{split}$ (3)
@@
The perturbation $\delta$ is constrained within an $\epsilon$-ball to balance attack effectiveness and imperceptibility, ensuring the $I+\delta$ remains visually indistinguishable from $I$. We minimize the $\ell_{\infty}$-norm distance between normalized feature representations to ensure scale-invariant alignment in the visual latent space, enhancing robustness of optimization process across different vision encoders.
@@
Specifically, we optimize Eq. (3) using the SSA-CWA Chen et al. (2024) algorithm, which iteratively computes gradients across the ensemble of surrogate models to produce the visual perturbation $\delta$, thereby generating adversarial images embedded with malicious semantics.
@@ 4.2 Textual Guidance Enhancement
While the Visual Latent Alignment manipulates the agent’s visual input to prioritize malicious instructions, it alone may not guarantee precise control over the agent’s output. To overcome this, we inject malicious content into the textual command input $C$ using injection function $\eta(\cdot)$ defined in Sec. 3.1, converting it into a deceptive command $C^{\prime}$. While the visual injection embeds malicious semantics through the input end, textual injection steers the agent core planners’ response towards compliance with attacker’s request from output end, forming a complementary bidirectional attack strategy. The manipulated textual command works synergistically with visual injection, fully exploiting vulnerabilities in both input processing and output generation pipeline of multimodal agents.
@@
Specifically, we initialize $C$ from random string and iteratively optimize it to maximize the likelihood of generating malicious action $a^{*}$ defined in Eq. (2), inducing the agent to generate output aligned with attacker’s adversarial goal. However, directly optimizing this objective with respect to $C$ under black-box setting is challenging, as multimodal agents typically have high-dimensional inputs and complex multimodal data processing pipelines, making end-to-end gradient computation infeasible Zhang et al. (2025). To circumvent this, we adopt the transfer attack strategy, leveraging open-source LLM as a surrogate model to approximate the victim agent’s inference process.
@@
To amplify the impact of the attack, we construct the defense-aware system prompt through role-driven analysis and optimize the malicious command based on it. The constructed system prompt with reinforced defense against prompt injection may push textual optimization to discover the command $C^{\prime}$ that overcomes such strong protection. This adversarially leads to the malicious command $C^{\prime}$ with enhanced control over the victim agent’s output. Specifically, we apply adversarial meta prompting Zhang et al. (2025) method to generate the construction prompts for different agents with powerful language model (i.e., GPT-4 OpenAI (2023)). Meta prompt is a rule-based functorial prompt template, which can be used to efficiently guide LLM to construct defensive system prompts for agents with diverse roles. Let $R$ be denoted as the role description of a victim agent, which is known to the attacker in our threat model, and let $T(\cdot,\cdot)$ represent the defense-aware meta-prompting template, the loss function for textual injection, $\mathcal{L}_{t}$, is defined as:
$\mathcal{L}_{t}(C)=-\log p(a^{*}\mid\mathcal{M}(T(R,\ r)),\ d,\ C),$ (4)
where $\mathcal{M}$ indicates the LLM for system prompt generation, and $r$ denotes the defensive rule against prompt injection. We minimize the negative log likelihood of the target action $a^{*}$ to maximize the probability of the agent generating response aligned with the malicious instruction $d$. To optimize Eq. (4), we employ the GCG algorithm, which iteratively computes the gradient of $\mathcal{L}_{t}$ with respect to tokens in $C$ and generates the optimal deceptive command. It has proven to be effective in transfer-based attacks Zou et al. (2023).